OpenStudy (anonymous):

Factor completely. 6x²+ 3x - 63 A. (3x + 9)(2x - 7) B. 3(2x²+ x - 21) C. (6x + 21)(x - 3) D. 3(2x + 7)(x - 3)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

just use mathway.com

OpenStudy (anonymous):

but it dont let me paste

OpenStudy (anonymous):

just type it lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

- 3x + 7y = 2 - 2x - 3y = - 4

OpenStudy (anonymous):

what are those equations for